Hello i just need to know how to set this question up. Like how would it go in the box? thanks.. A druggist mixed 70 cc (cubic centimeters) of a 25% alcohol solution with 40 cc of pure alcohol. Find the percent concentration of the resulting mixture. Round to the nearest tenth of a percent.
Answer by richwmiller(17219)

.25*70+1*40=110*x
17.5+40=110x
57.5=110x
57.5/110=x
.523=x
x=52.3%